# Is Epistemic Uncertainty Faithfully Represented by Evidential Deep Learning Methods?

### v2

@@ 2 Problem definition and literature review
Consider a classification or regression task with instance space $\mathcal{X}$ and label space $\mathcal{Y}$. Without loss of generality, we assume $\mathcal{X}=\mathbb{R}^{D}$ with $D$ the dimensionality of the data. Similarly, we assume $\mathcal{Y}=\mathbb{R}^{D^{\prime}}$ in the case of regression, $\mathcal{Y}=\{1,\ldots,K\}$ in $K$-class classification, and $\mathcal{Y}=\mathbb{N}$ in the case of counts. Furthermore, we have training data of the form $\mathcal{D}=\big{\{}\big{(}\boldsymbol{x}_{i},y_{i}\big{)}\big{\}}_{i=1}^{N}$, in which each $(\boldsymbol{x}_{i},y_{i})$ is i.i.d. sampled via a joint probability measure $P$ on $\mathcal{X}\times\mathcal{Y}$. Correspondingly, each feature vector $\boldsymbol{x}\in\mathcal{X}$ is associated with a conditional distribution $p(\cdot\,|\,\boldsymbol{x})$ on $\mathcal{Y}$, such that $p(y\,|\,\boldsymbol{x})$ is the probability to observe label $y$ as an outcome given $\boldsymbol{x}$. For the following, we denote $\mathbb{P}(A)$ as the set of all probability distributions on the set $A$.
@@ 2.1 First-order risk minimization
The classical statistics and machine learning literature usually considers first-order distributions during risk minimization. More specifically, one has a hypothesis space $\mathcal{H}_{1}$ containing $\Theta\rightarrow\mathbb{P}(\mathcal{Y})$ mappings with $\Theta$ a parameter space. In fact, $\mathcal{H}_{1}$ defines probability density (or mass) functions $p(y\,|\,\boldsymbol{\theta})$ from a parametric family with parameter vector $\boldsymbol{\theta}\in\Theta$, for continuous (or discrete) random variables. In a supervised learning context, $\boldsymbol{\theta}$ is a function of the features: we will use the notation $\boldsymbol{\theta}(\boldsymbol{x};\boldsymbol{\Phi})$, being an $\mathcal{X}\rightarrow\Theta$ function that is parameterized by a set of parameters $\boldsymbol{\Phi}$. Risk minimization is then usually defined as:
@@ 2.3 The exponential family
In regression, $\mathcal{H}_{1}$ typically represents the set of Gaussian distributions with parameters $\boldsymbol{\theta}=(\mu,\sigma^{2})$. $\mathcal{H}_{2}$ contains all normal-inverse Gamma (NIG) distributions with parameters $\boldsymbol{m}=(\gamma,\nu,\alpha,\beta)$. Second-order risk minimization is typically realized via a neural network architecture with one or multiple feature learning layers, represented by a mapping $\boldsymbol{\Psi}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{C}$, followed by four output neurons with functions $\gamma(\boldsymbol{x};\boldsymbol{\Psi})=\boldsymbol{w}_{1}^{\top}\boldsymbol{\Psi}(\boldsymbol{x})$, $\nu(\boldsymbol{x};\boldsymbol{\Psi})=\exp\left\{\boldsymbol{w}_{2}^{\top}\boldsymbol{\Psi}(\boldsymbol{x})\right\}$, $\alpha(\boldsymbol{x};\boldsymbol{\Psi})=\exp\left\{\boldsymbol{w}_{3}^{\top}\boldsymbol{\Psi}(\boldsymbol{x})\right\}$ and, $\beta(\boldsymbol{x};\boldsymbol{\Psi})=\exp\left\{\boldsymbol{w}_{4}^{\top}\boldsymbol{\Psi}(\boldsymbol{x})\right\}$. Here the exponential function guarantees that the parameters $\nu$, $\alpha$ and $\beta$ are positive. Sometimes the softplus function is used as an alternative (Amini et al., 2020). The parameter set $\boldsymbol{\Phi}$ is hence given by $\{\boldsymbol{w}_{1},\boldsymbol{w}_{2},\boldsymbol{w}_{3},\boldsymbol{w}_{4}\}$ and all the parameters considered in the feature learning layers.
@@
Figure 1: General overview of a neural network architecture for first-order (left) and second-order (right) risk minimization. While the first-order model learns the Bernoulli parameter $\theta$ of the data generating distribution directly, the second-order model predicts the parameters $\alpha$ and $\beta$ of a Beta distribution. The example is aligned with our experimental setup, where the feature space is one-dimensional.
@@ 2.4 Quantifying epistemic uncertainty
$I(\boldsymbol{m})=\mathrm{H}(\mathbb{E}_{\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x}_{i};\boldsymbol{\Phi}))}\!\left[p(y\,|\,\boldsymbol{\theta})\right])-\mathbb{E}_{\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x}_{i};\boldsymbol{\Phi}))}\!\left[\mathrm{H}(p(y\,|\,\boldsymbol{\theta})\right])\,,$
@@
An alternative scalar quantity for epistemic uncertainty is the entropy of $p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x}_{i};\boldsymbol{\Phi}))$. As far as we know, this measure has only been considered in one recent paper (Kotelevskii et al., 2023), which is surprising, because it is often used as a regularizer– see Section 3.4 for a more in-depth discussion. This last paper is also the only paper that consistently chooses the same measure for regularization and quantification of epistemic uncertainty. Roughly speaking, one can say that entropy analyzes the uncertainty of $\boldsymbol{\theta}$ under $p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x}_{i};\boldsymbol{\Phi}))$, whereas the covariance matrix rather quantifies the spread around the mean of $\boldsymbol{\theta}$. In Appendix A we provide entropy formulas for NIG, Dirichlet and Gamma distributions.
@@ 3.1 Reference second-order distribution
The question regarding the “right” representation of epistemic uncertainty is a difficult one, akin to the fundamental question of how to do statistical inference in the right way Martin (2023). Needless to say, a definite answer to this question cannot be expected. For the purpose of our analysis, we opt for a pragmatic approach specifically tailored to our setting, in which the second-order model is essentially meant to capture the uncertainty of a first-order predictor learned from the training data. The latter is the empirical minimizer of (1), which is extended to (2) or (3) by the second-order learner. The randomness of this minimizer is due to the randomness of the training data $\mathcal{D}$ [^1] In practice, additional randomness might be caused by the learning algorithm, which might not be guaranteed to deliver the true minimizer. Theoretically, the uncertainty is therefore reflected by the distribution of the first-order predictor induced by the sampling process that generates the training data. This leads us to the notion of a reference distribution, and the idea that the epistemic uncertainty predicted by the second-order learner should ideally be represented by this distribution.
@@
Let $\boldsymbol{\theta}_{\mathcal{D}_{N}}(\boldsymbol{x};\boldsymbol{\Phi}_{\mathcal{D}_{N}})$ denote the minimizer of (1) for a training set $\mathcal{D}_{N}$ of size $N$, where $\mathcal{D}_{N}\sim P^{N}$. For a given $\boldsymbol{x}\in\mathcal{X}$, we define the reference second-order distribution as
@@
Obviously, the reference distribution cannot be computed, because the true data distribution $P$ is unknown. Yet, it can be approximated in simulation studies, a technique that we shall use in our experiments later on. Moreover, note that the reference second-order distribution depends on various design choices for first-order risk minimization, such as the hypothesis space, loss function and optimization algorithm. As such, we believe that the reference should not necessarily be interpreted as a ground-truth epistemic uncertainty representation, but it defines the distribution that might be expected when compatible hypothesis spaces, loss functions and optimization algorithms are used for first-order and second-order risk minimization (i.e. for second-order risk minimization one could use the same neural network architecture and optimizer as for first-order risk minimization, and the likelihood of the former might correspond to the conjugate prior of the latter). The second-order prediction $p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x};\boldsymbol{\Phi}))$ should ideally be close to the reference $q_{N}(\boldsymbol{\theta}\,|\,\boldsymbol{x})$, which can be quantified with any distance metric on probability distributions (Rachev et al., 2013).
@@ 3.2 Unregularized inner loss minimization
$p(y\,|\,\boldsymbol{m}(\boldsymbol{x}))=\mathbb{E}_{\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x};\boldsymbol{\Phi}))}\,p(y\,|\,\boldsymbol{\theta}).$
@@
$\hat{\boldsymbol{\Phi}}=\operatorname{arg}\min_{\boldsymbol{\Phi}}\,\sum_{i=1}^{N}L_{1}(y_{i},p(y_{i}\,|\,\boldsymbol{m}(\boldsymbol{x}_{i};\boldsymbol{\Phi})))\,.$ (5)
@@ 3.3 Unregularized outer loss minimization
$\hat{\boldsymbol{\Phi}}=\operatorname{arg}\min_{\boldsymbol{\Phi}}\sum_{i=1}^{N}\mathbb{E}_{\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x}_{i};\boldsymbol{\Phi}))}L_{1}(y_{i},p(y\,|\,\boldsymbol{\theta}))\,.$ (6)
Similarly as for inner loss minimization, closed-form expressions exist for the second-order loss (see Appendix D and (Charpentier et al., 2022)). Yet, it turns out that optimization problem (6) has different characteristics than optimization problem (5). Here comes our main result for this subsection.
@@ 3.4 Regularized risk minimization
The proof (see Appendix F) is based on an equivalent formulation of (2) and (3) as constrained optimization problems. This reformulation reveals the role of the regularizer: it penalizes predictions with a low epistemic uncertainty. More specifically, the epistemic uncertainty for all training data points is summed up, and this sum is fixed to a certain level determined by $\lambda$. In other words, the constraint in the two optimization problems defines an epistemic uncertainty budget that cannot be exceeded for the training data points. Now, in light of Def. 3.1 we consider for all $\boldsymbol{x}\in\mathcal{X}$ a probability distribution defined over the empirical risk minimizer $\hat{\boldsymbol{\theta}}$ and epistemic uncertainty quantified by its entropy. If $\lambda$ is chosen sufficiently large, the allowed epistemic uncertainty budget will be too small to represent the epistemic uncertainty of the reference distribution, expressed by entropy or any other epistemic uncertainty measure.
@@ 4 Experimental analysis
The main goal of the experiments is to illustrate that our theoretical observations also arise in practice. As argued before, the faithfulness of an epistemic uncertainty representation w.r.t. Def. 3.1 can only be empirically verified via simulation studies that assume access to the reference distribution $P$. We performed two simulation studies, covering the classification and the regression case. For the sake of brevity, we only describe the classification case here, but the results for the regression case also confirm the theoretical results (see Appendix G).
@@ 4.1 Setup
We consider a binary classification setup, because the reference distribution and the second-order fit can be easily visualized in that case. We generate synthetic training data $\mathcal{D}_{N}=\{(x_{i},y_{i})\}_{i=1}^{N}$ of size $N$, with $x_{i}\sim\mathrm{Unif}([0,0.5])$, and $y_{i}\sim\mathrm{Bern}\big{(}\theta(x_{i})\big{)}$. The Bernoulli parameter $\theta$ is described as a rather complicated function of the one-dimensional feature $x$: $\theta(x_{i})=0.5\cdot(0.8\cdot\sin(2\pi x_{i})+1)$, thus $\theta(x_{i})\in[0.1,0.9]$. Intuitively, this looks like a very simple problem, but the aleatoric uncertainty is very high for most $x$, so one needs a lot of training data to estimate $\theta(x)$ with high precision.
@@
To train the first-order model in (1), negative log-likelihood is used as a loss function (i.e., logistic loss). The second-order models predict the parameters $\boldsymbol{m}(x)=(\alpha(x),\beta(x))^{\top}$ of a Beta distribution and are trained using the loss functions derived from outer and inner expectation minimization, respectively. In both cases, we use closed-form expressions for the loss functions, which avoids the need for sampling– see Appendices B and D for derivations. For the regularized models, we use negative entropy of the predicted distribution with $\lambda=0.01$.

### v3

@@ 2 Problem definition and literature review
Consider a classification or regression task with instance space $\mathcal{X}$ and label space $\mathcal{Y}$. Without loss of generality, we assume $\mathcal{X}=\mathbb{R}^{D}$ with $D$ the dimensionality of the data. Similarly, we assume $\mathcal{Y}=\mathbb{R}^{D^{\prime}}$ in the case of regression, $\mathcal{Y}=\{1,\ldots,K\}$ in $K$-class classification, and $\mathcal{Y}=\mathbb{N}$ in the case of counts. Furthermore, we have training data of the form $\mathcal{D}=\big{\{}\big{(}\boldsymbol{x}_{i},y_{i}\big{)}\big{\}}_{i=1}^{N}$, in which each $(\boldsymbol{x}_{i},y_{i})$ is i.i.d. sampled via a joint probability measure $P$ on $\mathcal{X}\times\mathcal{Y}$. Correspondingly, each feature vector $\boldsymbol{x}\in\mathcal{X}$ is associated with a conditional distribution $p(\cdot\,|\,\boldsymbol{x})$ on $\mathcal{Y}$, such that $p(y\,|\,\boldsymbol{x})$ is the probability to observe label $y$ as an outcome given $\boldsymbol{x}$. More generally, $\mathbb{P}(A)$ will denote the set of all probability distributions on the set $A$.
@@ 2.1 First-order risk minimization
The classical statistics and machine learning literature usually considers first-order distributions during risk minimization. More specifically, one has a hypothesis space $\mathcal{H}_{1}$ containing $\Theta\rightarrow\mathbb{P}(\mathcal{Y})$ mappings with $\Theta$ a parameter space. In fact, $\mathcal{H}_{1}$ defines probability density (or mass) functions $p(y\,|\,\boldsymbol{\theta})$ from a parametric family with parameter vector $\boldsymbol{\theta}\in\Theta$, for continuous (or discrete) random variables. In a supervised learning context, $\boldsymbol{\theta}$ is a function of the features: we will use the notation $\boldsymbol{\theta}(\boldsymbol{x};\boldsymbol{\Phi})$, being an $\mathcal{X}\rightarrow\Theta$ function that is parameterized by a set of parameters $\boldsymbol{\Phi}$. Risk minimization is then usually defined as
@@ 2.3 The exponential family
In regression, $\mathcal{H}_{1}$ typically represents the set of Gaussian distributions with parameters $\boldsymbol{\theta}=(\mu,\sigma^{2})$. $\mathcal{H}_{2}$ contains all normal-inverse Gamma (NIG) distributions with parameters $\boldsymbol{m}=(\gamma,\nu,\alpha,\beta)$. Second-order risk minimization is typically realized via a neural network architecture with one or multiple feature learning layers, represented by a mapping $\boldsymbol{\Psi}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{C}$, followed by four output neurons with functions $\gamma(\boldsymbol{x};\boldsymbol{\Psi})=\boldsymbol{w}_{1}^{\top}\boldsymbol{\Psi}(\boldsymbol{x})$, $\nu(\boldsymbol{x};\boldsymbol{\Psi})=\exp\!\left\{\boldsymbol{w}_{2}^{\top}\boldsymbol{\Psi}(\boldsymbol{x})\!\right\}$, $\alpha(\boldsymbol{x};\boldsymbol{\Psi})=\exp\!\left\{\boldsymbol{w}_{3}^{\top}\boldsymbol{\Psi}(\boldsymbol{x})\!\right\}$ and, $\beta(\boldsymbol{x};\boldsymbol{\Psi})=\exp\!\left\{\boldsymbol{w}_{4}^{\top}\boldsymbol{\Psi}(\boldsymbol{x})\!\right\}$. Here the exponential function guarantees that the parameters $\nu$, $\alpha$ and $\beta$ are positive. Sometimes the softplus function is used as an alternative (Amini et al., 2020). The parameter set $\boldsymbol{\Phi}$ is hence given by $\{\boldsymbol{w}_{1},\boldsymbol{w}_{2},\boldsymbol{w}_{3},\boldsymbol{w}_{4}\}$ and all the parameters of the feature learning layers.
@@
Figure 1: General overview of a neural network architecture for first-order (left) and second-order (right) risk minimization. While the first-order model learns the Bernoulli parameter $\theta$ of the data generating distribution directly, the second-order model predicts the parameters $\boldsymbol{m}=(\alpha,\beta)^{\top}$ of a Beta distribution, which defines a distribution with density $f(\theta|x,\alpha,\beta)=\frac{1}{B(\alpha,\beta)}\theta^{\alpha-1}(1-\theta)^{\beta-1}$ over $\theta$. The example is aligned with our experimental setup, where the feature space is one-dimensional.
@@ 2.4 Quantifying epistemic uncertainty
$I(\boldsymbol{m})=\mathrm{H}(\mathbb{E}_{\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x}_{i};\boldsymbol{\Phi}))}\!\left[p(y\,|\,\boldsymbol{\theta})\right])\\ -\mathbb{E}_{\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x}_{i};\boldsymbol{\Phi}))}\!\left[\mathrm{H}(p(y\,|\,\boldsymbol{\theta})\right])\,,$
@@
An alternative scalar quantity for epistemic uncertainty is the entropy of $p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x}_{i};\boldsymbol{\Phi}))$. As far as we know, this measure has only been considered in one recent paper (Kotelevskii et al., 2023), which is surprising, because it is often used as a regularizer — see Section 3.4 for a more in-depth discussion. This last paper is also the only paper that consistently chooses the same measure for regularization and quantification of epistemic uncertainty. Roughly speaking, one can say that entropy analyzes the uncertainty of $\boldsymbol{\theta}$ under $p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x}_{i};\boldsymbol{\Phi}))$, whereas the covariance matrix rather quantifies the spread around the mean of $\boldsymbol{\theta}$. In Appendix A we provide entropy formulas for NIG, Dirichlet and Gamma distributions.
@@ 3.1 Reference second-order distribution
The question regarding the “right” representation of epistemic uncertainty is a difficult one, akin to the fundamental question of how to do statistical inference in the right way (Martin, 2023). Needless to say, a definite answer to this question cannot be expected. For the purpose of our analysis, we opt for a pragmatic approach specifically tailored to our setting, in which the second-order model is essentially meant to capture the uncertainty of a first-order predictor learned from the training data. The latter is the empirical minimizer of (1), which is extended to (2) or (3) by the second-order learner. The randomness of this minimizer is due to the randomness of the training data $\mathcal{D}$ [^1] In practice, additional randomness might be caused by the learning algorithm, which might not be guaranteed to deliver the true minimizer. Theoretically, the uncertainty is therefore reflected by the distribution of the first-order predictor induced by the sampling process that generates the training data. This leads us to the notion of a reference distribution, and the idea that the epistemic uncertainty predicted by the second-order learner should ideally be represented by this distribution.
@@
Let $\boldsymbol{\theta}_{\mathcal{D}_{N}}(\boldsymbol{x};\boldsymbol{\Phi}_{\mathcal{D}_{N}})$ denote the minimizer of (1) for a training set $\mathcal{D}_{N}$ of size $N$, where $\mathcal{D}_{N}\sim P^{N}$. For a given $\boldsymbol{x}\in\mathcal{X}$, we define the reference second-order distribution as the conditional distribution induced by $\boldsymbol{\theta}_{\mathcal{D}_{N}}(\boldsymbol{x};\boldsymbol{\Phi}_{\mathcal{D}_{N}})$:
@@
Intuitively, this definition follows a frequentist statistics viewpoint, in which a ground-truth model is estimated using limited training data, and epistemic uncertainty is caused by the sum of squared bias and variance of the estimate. The bias-variance decomposition is frequentist in nature, because it assumes a ground-truth model that needs to be estimated using a training dataset of limited size. Typically, the assumption is made that a training dataset of a given size can be resampled to quantify bias (by comparison of the mean to the ground truth) and variance in an objective way. More precisely, one can resample the training datasets $d$ times from the underlying distribution, resulting in $d$ empirical risk minimisers $\{\hat{\boldsymbol{\theta}}_{1},\dots,\hat{\boldsymbol{\theta}}_{d}\}$, which then allow to compute the empirical (cumulative) distribution function:
$\hat{Q}_{N}(\boldsymbol{\theta}|\boldsymbol{x}):=\frac{1}{d}\sum_{i=1}^{d}\mathbb{I}(\hat{\boldsymbol{\theta}_{i}}(\boldsymbol{x},\boldsymbol{\Phi})\leq\boldsymbol{\theta}).$
Moreover, note that the reference second-order distribution depends on various design choices for first-order risk minimization, such as the hypothesis space, loss function and optimization algorithm. As such, we believe that the reference should not necessarily be interpreted as a ground-truth epistemic uncertainty representation, but it defines the distribution that might be expected when compatible hypothesis spaces, loss functions and optimization algorithms are used for first-order and second-order risk minimization (i.e. for second-order risk minimization one could use the same neural network architecture and optimizer as for first-order risk minimization, and the likelihood of the former might correspond to the conjugate prior of the latter). The second-order prediction $p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x};\boldsymbol{\Phi}))$ should ideally be close to the reference $q_{N}(\boldsymbol{\theta}\,|\,\boldsymbol{x})$, which can be quantified with any distance metric on probability distributions (Rachev et al., 2013).
@@ 3.2 Unregularized inner loss minimization
$p(y\,|\,\boldsymbol{m}(\boldsymbol{x}))=\mathbb{E}_{\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x};\boldsymbol{\Phi}))}\!\left[p(y\,|\,\boldsymbol{\theta})\right].$
@@
$\hat{\boldsymbol{\Phi}}=\operatorname{arg}\min_{\boldsymbol{\Phi}}\,\sum_{i=1}^{N}L_{1}(y_{i},p(y\,|\,\boldsymbol{m}(\boldsymbol{x}_{i};\boldsymbol{\Phi})))\,.$ (5)
@@ 3.3 Unregularized outer loss minimization
$\hat{\boldsymbol{\Phi}}=\operatorname{arg}\min_{\boldsymbol{\Phi}}\sum_{i=1}^{N}\mathbb{E}_{\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\,|\,\boldsymbol{m}(\boldsymbol{x}_{i};\boldsymbol{\Phi}))}\!\left[L_{1}(y_{i},p(y\,|\,\boldsymbol{\theta}))\right]\,.$ (6)
Similarly as for inner loss minimization, closed-form expressions exist for the second-order loss (see Appendix D and Charpentier et al. (2022)). Yet, it turns out that optimization problem (6) has different characteristics than optimization problem (5). Here comes our main result for this subsection:
@@
Figure 3: Empirical Wasserstein-$1$ distance between the reference and the estimated second-order distributions for $\lambda\in[0.0,0.001,0.01,0.05,0.1,0.5]$. The distance is calculated for $100$ equidistant points in $\mathcal{X}=[0,1]$, by evaluating empirical and second-order distributions in the parameter space and calculating the average $L_{1}$ distance.
@@ 3.4 Regularized risk minimization
The proof (see Appendix F) is based on an equivalent formulation of (2) and (3) as constrained optimization problems. This reformulation reveals the role of the regularizer: it penalizes predictions with a low epistemic uncertainty. More specifically, the estimated epistemic uncertainty for all training data points is summed up, and this sum is fixed to a certain level determined by $\lambda$. In other words, the constraint in the two optimization problems defines an epistemic uncertainty budget that cannot be exceeded for the training data points. Now, in light of Def. 3.1 we consider for all $\boldsymbol{x}\in\mathcal{X}$ a probability distribution defined over the empirical risk minimizer $\hat{\boldsymbol{\theta}}$ and epistemic uncertainty quantified by its entropy. If $\lambda$ is chosen sufficiently small, the allowed epistemic uncertainty budget will be too small to represent the epistemic uncertainty of the reference distribution, expressed by entropy or any other epistemic uncertainty measure.
@@ 4 Experimental analysis
The main goal of the experiments is to illustrate that our theoretical observations also arise in practice. As argued before, the faithfulness of an epistemic uncertainty representation w.r.t. Def. 3.1 can only be empirically verified via simulation studies that assume access to the reference distribution $P$. We performed two simulation studies, covering the classification and the regression case. For the sake of brevity, we only describe the classification case here, but the results for the regression case also confirm the theoretical results (see Appendix G). The code for reproducing the experiments is available on GitHub. [^2] https://github.com/mkjuergens/EpistemicUncertaintyAnalysis
@@ 4.1 Setup
We consider a binary classification setup, because the reference distribution and the second-order fit can be easily visualized in that case. We generate synthetic training data $\mathcal{D}_{N}=\{(x_{i},y_{i})\}_{i=1}^{N}$ of size $N$, with $x_{i}\sim\mathrm{Unif}([0,0.5])$, and $y_{i}\sim\mathrm{Bern}\big{(}\theta(x_{i})\big{)}$. The Bernoulli parameter $\theta$ is described as a function of the one-dimensional features $x_{i}$: $\theta(x_{i})=0.5+0.4\,\sin 2\pi x_{i}$, thus $\theta(x_{i})\in[0.5,0.9]$. Intuitively, this looks like a very simple problem, but the aleatoric uncertainty is very high for most $x$, so one needs a lot of training data to estimate $\theta(x)$ with high precision.
@@
To train the first-order model in (1), negative log-likelihood is used as a loss function (i.e., logistic loss). The second-order models predict the parameters $\boldsymbol{m}(x)=(\alpha(x),\beta(x))^{\top}$ of a Beta distribution and are trained using the loss functions derived from outer and inner expectation minimization, respectively. In both cases, we use closed-form expressions for the loss functions, which avoids the need for sampling—see Appendices B and D for derivations. For the regularized models, we use negative entropy of the predicted distribution with $\lambda=0.01$.